# FLLIC: Functionally Lossless Image Compression

### v3

@@ I Introduction
We provide a visual comparison between the traditional frameworks for noisy image compression and the proposed functionally lossless compression method in Fig. 1. In the current practice, a noisy image is either directly losslessly compressed or first denoised and then losslessly compressed. These two approaches are both sub-optimal in terms of rate-distortion metric. Direct lossless compression, by struggling to preserve noises, suffers in both aspects of fidelity and bit rate. It is detrimental to the transmission and the subsequent machine vision tasks. The cascaded approach of denoising followed by lossless compression is complex and computationally expensive. In contrast, the proposed functionally lossless image compression (FLLIC) method jointly optimizes denoising and compression, and achieves higher coding efficiency and lower latency.
@@ II-B Lossless and Near-lossless Image Compression
Several representative works have been proposed in this area: Zhou et al[55, 56] introduced a soft-decoding approach to reduce the $\ell_{2}$ distortion of $\ell_{\infty}$-decoded images. Their method formulates soft decoding as an image restoration problem, leveraging the tight error bounds provided by $\ell_{\infty}$-constrained coding and employing a context-based modeler to handle quantization errors. Chuah et al[57] replaced the traditional uniform scalar quantizer used in near-lossless image coding with a set of context-based, $\ell_{2}$-optimized quantizers. Their approach minimizes a weighted sum of $\ell_{2}$ distortion and entropy while maintaining a strict $\ell_{\infty}$ error bound. Li et al[58] proposed a sparsity-driven restoration technique to improve the coding performance of $\ell_{\infty}$-decoded images. Their method exploits the sparsity characteristics of the image for efficient restoration. Florea et al[59] focused on optimizing mesh codecs with respect to the $\ell_{\infty}$ metric. They proposed novel data-dependent formulations for $\ell_{\infty}$ distortion and incorporated $\ell_{\infty}$-based estimators into a state-of-the-art wavelet-based semi-regular mesh codec. Zhang et al[25] developed an asymmetric image compression system that features lightweight $\ell_{\infty}$-constrained predictive encoding combined with a computationally intensive, learning-based soft decoding process to achieve high-fidelity reconstruction. Bai et al[60, 61] proposed a joint framework for lossy image compression and residual compression to achieve $\ell_{\infty}$-constrained near-lossless compression. Specifically, their approach first obtains a lossy reconstruction of the input image using lossy compression and then uniformly quantizes the residual to satisfy a tight $\ell_{\infty}$ error bound. These works collectively demonstrate the advancements in near-lossless image compression, focusing on balancing rate-distortion performance while strictly adhering to the $\ell_{\infty}$ error constraint.
@@ II-C Joint Image Compression and Denoising
Image compression and image denoising have been thoroughly studied by researchers in both camps of traditional image processing and modern deep learning. However, the joint image compression and denoising task has been little explored. Very few papers addressed this topic. Testolina et al.[62] investigated the integration of denoising convolutional layers in the decoder of a learning-based compression network. Ranjbar et al.[63] presented a learning-based image compression framework where image denoising and compression are performed jointly. The latent space of the image codec is organized in a scalable manner such that the clean image can be decoded from a subset of the latent space, while the noisy image is decoded from the full latent space at a higher rate. Cheng et al.[64] proposed to optimize the image compression algorithm to be noise-aware as joint denoising and compression. The key is to transform the original noisy images to noise-free bits by eliminating the undesired noise during compression, where the bits are later decompressed as clean images. Huang et al.[65] proposed an efficient end-to-end image compression network, named Noise-Adaptive ResNet VAE (NARV), aiming to handle both clean and noisy input images of different noise levels in a single noise-adaptive compression network without adding nontrivial processing time. Although these works realized the significance of the joint image compression and denoising problem, they just combined image denoising with lossy compression task, with no regard to the lossless compression problem. To our best knowledge, we are the first to investigate the joint image denoising and lossless compression problem.
@@ III-B Theoretical Analysis of Clean Entropy Estimation
Let $X\in\mathbb{R}^{n}$ be a random vector representing the clean image, and $Y=X+N$ its noisy observation, where $N$ is an additive noise vector independent of $X$. We assume both $X$ and $N$ are zero-mean Gaussian vectors. Although real images are not strictly Gaussian, Gaussian modeling is a common and analytically tractable approximation that often provides valuable insight. We denote the covariance matrices as $\Sigma_{X}=\mathbb{E}[XX^{\top}]$ and $\Sigma_{N}=\mathbb{E}[NN^{\top}]$. Without loss of generality, let $\Sigma_{N}=\sigma_{N}^{2}I$, i.e., isotropic Gaussian noise with variance $\sigma_{N}^{2}$. Since $Y=X+N$, it follows that $\Sigma_{Y}=\Sigma_{X}+\Sigma_{N}$.
@@
This relationship links the entropy gap to the signal-to-noise ratio (SNR) characteristics of $X$. Eigenvalues $\lambda_{i}$ represent the variance of $X$ along principal components. Signal components with smaller $\lambda_{i}$ are more easily overwhelmed by noise, thus increasing $h(Y)$ relative to $h(X)$. Building on this understanding, we can derive the following key insights and analyses:
@@ III-C Practical Estimation of Clean Image Entropy
While the theoretical framework offers a clear understanding of $H(I)$, direct estimation in practical scenarios remains challenging. We address this by introducing a deep neural network (DNN) designed to approximate $H(I)$ from the noisy observation $I_{n}$, an estimate of $H(I_{n})$, and (if available) the noise variance $\sigma_{N}^{2}$. This approach leverages both the distributional properties of the noisy data and any accessible entropy information about $I_{n}$ to infer $H(I)$ without direct access to the clean image $I$ during training.
Formally, let $I\in\mathbb{R}^{H\times W\times C}$ be the clean image, and $N\sim\mathcal{N}(0,\sigma_{N}^{2}I)$ an additive Gaussian noise field, giving $I_{n}=I+N$. We assume:
@@
where $\hat{H}(I_{n})$ is an estimate obtained by applying a mathematically lossless image compression (MLLIC) model to $I_{n}$. Our goal is to construct an estimator:
@@
where $f_{\phi}(\cdot)$ is a DNN with learnable parameters $\phi$. This network is trained to approximate the mapping from $(I_{n},\hat{H}(I_{n}),\sigma_{N})$ to $H(I)$.
To implement $f_{\phi}(\cdot)$, we employ a two-stage approach consisting of a feature extraction module followed by a regression module. First, a network $g_{\theta}(\cdot)$ with parameters $\theta$ is used to extract high-level features $F(I_{n})=g_{\theta}(I_{n})$ from the noisy input $I_{n}$:
@@
Specifically, we adopt a cascade of residual blocks (ResBlocks) and downsampling residual blocks (Down-ResBlocks). Each ResBlock is defined as:
$F_{\text{res}}=F_{\text{in}}+h_{\text{res}}(F_{\text{in}}),$ (12)
where $F_{\text{in}}$ is the input feature map to the block and $h_{\text{res}}(\cdot)$ is a nonlinear transformation composed of convolutional and activation layers. Each Down-ResBlock is similarly defined, but incorporates downsampling to progressively reduce the spatial resolution:
$F_{\text{down}}=d_{\text{res}}(F_{\text{in}})+F_{\text{in}},$ (13)
where $d_{\text{res}}(\cdot)$ includes strided convolutions or pooling layers to downsample the feature maps, alongside residual connections to stabilize training.
By stacking multiple ResBlocks and Down-ResBlocks, the network $g_{\theta}(\cdot)$ encodes $I_{n}$ into a compact and informative representation. After passing through a sequence of these blocks, we obtain a final feature representation $F(I_{n})$ that captures both global and local structures, as well as noise characteristics.
Once $F(I_{n})$ is obtained, it is flattened into a vector $F_{v}=\mathrm{vec}(F(I_{n}))$. We then concatenate $F_{v}$, $\hat{H}(I_{n})$, and $\sigma_{N}$ into a single input vector $Z$ for the regression module:
$Z=[F_{v};\hat{H}(I_{n});\sigma_{N}].$ (14)
The regression module $r_{\psi}(\cdot)$, with parameters $\psi$, is a multilayer perceptron (MLP) that maps $Z$ to a scalar:
@@
To learn the parameters $\{\theta,\psi\}$, we use a dataset of image triplets $(I,I_{n},\sigma_{N})$ for which $H(I)$ is known or approximated via a high-fidelity reference. We minimize the following loss:
$\min_{\theta,\psi}\mathbb{E}\bigl{[}(\hat{H}(I)-H(I))^{2}\bigr{]}.$ (16)
In summary, this practical estimation framework employs a network of ResBlocks and Down-ResBlocks to extract a robust representation of the noisy observation $I_{n}$. The final regression module integrates these learned features with $\hat{H}(I_{n})$ and $\sigma_{N}$ to predict $\hat{H}(I)$. By jointly learning from noisy data and entropy estimates of $I_{n}$, the model provides a flexible and effective solution for estimating clean image entropy in real-world scenarios.
@@ III-D Content-Adaptive Quantization and Probability Estimation
The proposed framework integrates entropy-guided adaptive quantization and probability estimation to optimize joint denoising and compression. By incorporating clean image entropy $\hat{H}(I)$, latent features $\mathbf{y}$, and hyperlatents $\mathbf{z}$, the framework achieves precise control over quantization and accurate probability modeling, effectively balancing rate-distortion trade-offs.
@@
1. Quantization Map for Latent Features $\mathbf{y}$: The quantization step map $\mathbf{Q}_{\mathbf{y}}\in\mathbb{R}^{H\times W\times C}$ for the latent features $\mathbf{y}$ is computed as:
$\mathbf{Q}_{\mathbf{y}}=f_{\mathbf{Q_{y}}}(\mathbf{y},\hat{\mathbf{z}},\hat{H}(I)),$ (17)
where $\mathbf{y}\in\mathbb{R}^{H\times W\times C}$ represents the latent features, $\hat{\mathbf{z}}$ denotes the decoded hyperlatents, and $\hat{H}(I)$ is the estimated clean image entropy. The function $f_{\mathbf{Qy}}(\cdot)$ is a neural network parameterized to produce $\mathbf{Q}_{\mathbf{y}}$ by analyzing the spatial, channel-wise, and entropy-based characteristics of $\mathbf{y}$.
The quantization process for $\mathbf{y}$ is defined as:
$\mathbf{y}_{q}=\mathrm{round}\left(\frac{\mathbf{y}}{\mathbf{Q}_{\mathbf{y}}}\right),$ (18)
where division and rounding are performed element-wise. Regions with higher complexity or structural importance are assigned finer quantization steps, ensuring accurate reconstruction, while smoother regions are quantized more coarsely to save bits.
2. Quantization Map for Hyperlatents $\mathbf{z}$: The quantization step map $\mathbf{Q}_{\mathbf{z}}\in\mathbb{R}^{H^{\prime}\times W^{\prime}\times C^{\prime}}$ for the hyperlatents $\mathbf{z}$ is derived as:
$\mathbf{Q}_{\mathbf{z}}=f_{\mathbf{Q_{z}}}(\mathbf{z},\hat{H}(I)),$ (19)
where $\mathbf{z}\in\mathbb{R}^{H^{\prime}\times W^{\prime}\times C^{\prime}}$ represents the hyperlatents, and $f_{\mathbf{Qz}}(\cdot)$ is a neural network conditioned on $\mathbf{z}$ and $\hat{H}(I)$. This design ensures that the quantization map adapts to the global statistical properties of $\mathbf{z}$ and the entropy of the clean image.
The quantization process for $\mathbf{z}$ follows:
$\mathbf{z}_{q}=\mathrm{round}\left(\frac{\mathbf{z}}{\mathbf{Q}_{\mathbf{z}}}\right).$ (20)
This adaptive quantization efficiently encodes the hyperlatents while minimizing distortion.

### v4

@@ I Introduction
We provide a visual comparison between the traditional frameworks for noisy image compression and the proposed functionally lossless compression method in Figure 1. In the current practice, a noisy image is either directly losslessly compressed or first denoised and then losslessly compressed. These two approaches are both sub-optimal in terms of rate-distortion metric. Direct lossless compression, by struggling to preserve noises, suffers in both aspects of fidelity and bit rate. It is detrimental to the transmission and the subsequent machine vision tasks. The cascaded approach of denoising followed by lossless compression is complex and computationally expensive. In contrast, the proposed functionally lossless image compression (FLLIC) method jointly optimizes denoising and compression, and achieves higher coding efficiency and lower latency.
@@ II-B Lossless and Near-lossless Image Compression
Several representative works have been proposed in this area: Zhou et al[55, 56] introduced a soft-decoding approach to reduce the $\ell_{2}$ distortion of $\ell_{\infty}$-decoded images. Their method formulates soft decoding as an image restoration problem, leveraging the tight error bounds provided by $\ell_{\infty}$-constrained coding and employing a context-based modeler to handle quantization errors. Chuah et al[57] replaced the traditional uniform scalar quantizer used in near-lossless image coding with a set of context-based, $\ell_{2}$-optimized quantizers. Their approach minimizes a weighted sum of $\ell_{2}$ distortion and entropy while maintaining a strict $\ell_{\infty}$ error bound. Li et al[58] proposed a sparsity-driven restoration technique to improve the coding performance of $\ell_{\infty}$-decoded images. Their method exploits the sparsity characteristics of the image for efficient restoration. Florea et al[59] focused on optimizing mesh codecs with respect to the $\ell_{\infty}$ metric. They proposed novel data-dependent formulations for $\ell_{\infty}$ distortion and incorporated $\ell_{\infty}$-based estimators into a state-of-the-art wavelet-based semi-regular mesh codec. Zhang et al[25] developed an asymmetric image compression system that features lightweight $\ell_{\infty}$-constrained predictive encoding combined with a computationally intensive, learning-based soft decoding process to achieve high-fidelity reconstruction. Bai et al[60, 61] proposed a joint framework for lossy image compression and residual compression to achieve $\ell_{\infty}$-constrained near-lossless compression. Specifically, their approach first obtains a lossy reconstruction of the input image using lossy compression and then uniformly quantizes the residual to satisfy a tight $\ell_{\infty}$ error bound.
@@ II-C Joint Image Compression and Denoising
Image compression and image denoising have been thoroughly studied by researchers in both camps of traditional image processing and modern deep learning. However, the joint image compression and denoising task has been little explored. Very few papers addressed this topic. Testolina et al.[62] investigated the integration of denoising convolutional layers in the decoder of a learning-based compression network. Ranjbar et al.[63] presented a learning-based image compression framework where image denoising and compression are performed jointly. The latent space of the image codec is organized in a scalable manner such that the clean image can be decoded from a subset of the latent space, while the noisy image is decoded from the full latent space at a higher rate. Cheng et al.[64] proposed to optimize the image compression algorithm to be noise-aware as joint denoising and compression. The key is to transform the original noisy images to noise-free bits by eliminating the undesired noise during compression, where the bits are later decompressed as clean images. Huang et al.[65] proposed an efficient end-to-end image compression network, named Noise-Adaptive ResNet VAE (NARV), aiming to handle both clean and noisy input images of different noise levels in a single noise-adaptive compression network without adding nontrivial processing time. Recent work [66] has explored the joint design of denoising and compression, demonstrating that integrating the two tasks yields superior rate–distortion trade-offs compared to cascaded pipelines. While sharing this high-level motivation, our method fundamentally differs in both objective and formulation. Specifically, [66] focuses on lossy compression trained with classical rate–distortion objectives, whereas our work establishes the new paradigm of Functionally Lossless Image Compression (FLLIC), which aims at lossless coding of optimally denoised content to achieve zero degradation for downstream tasks.
Technically, FLLIC introduces entropy-guided, content-adaptive quantization driven by clean-entropy estimation and employs learned hyper-priors for precise entropy modeling. Moreover, we provide a theoretical analysis of the entropy gap between noisy and clean images to motivate our design—a perspective not covered in such prior works.
@@ III-B Theoretical Analysis of Clean Entropy Estimation
Gaussian assumption. Let $X\in\mathbb{R}^{n}$ be a random vector representing the clean image, and $Y=X+N$ its noisy observation, where $N$ is an additive noise vector independent of $X$. We assume both $X$ and $N$ are zero-mean Gaussian vectors. Although real images are not strictly Gaussian, Gaussian modeling is a common and analytically tractable approximation that often provides valuable insight. We denote the covariance matrices as $\Sigma_{X}=\mathbb{E}[XX^{\top}]$ and $\Sigma_{N}=\mathbb{E}[NN^{\top}]$. Without loss of generality, let $\Sigma_{N}=\sigma_{N}^{2}I$, i.e., isotropic Gaussian noise with variance $\sigma_{N}^{2}$. Since $Y=X+N$, it follows that $\Sigma_{Y}=\Sigma_{X}+\Sigma_{N}$.
@@
This relationship links the entropy gap to the signal-to-noise ratio (SNR) characteristics of $X$. Eigenvalues $\lambda_{i}$ represent the variance of $X$ along principal components. Signal components with smaller $\lambda_{i}$ are more easily overwhelmed by noise, thus increasing $h(Y)$ relative to $h(X)$.
Non-Gaussian extension. Now let $X$ be an arbitrary zero-mean vector with covariance $\Sigma_{X}$, still independent of the Gaussian noise $N\sim\mathcal{N}(0,\sigma_{N}^{2}I)$. Denote by $g_{Z}$ the Gaussian with the same mean and covariance as a random vector $Z$. It holds that
$D(Z\parallel g_{Z})=\frac{1}{2}\log\!\bigl{[}(2\pi e)^{n}\det(\Sigma_{Z})\bigr{]}-h(Z).$ (7)
Using this for both $X$ and $Y$, we get
$\begin{split}h(Y)-h(X)&=\frac{1}{2}\log\frac{\det(\Sigma_{X}+\sigma_{N}^{2}I)}{\det(\Sigma_{X})}\\[6.0pt] &\quad+\bigl{[}\,D(X\parallel g_{X})-D(Y\parallel g_{Y})\bigr{]}.\end{split}$ (8)
Because adding independent Gaussian noise makes a distribution more Gaussian,
$D(Y\parallel g_{Y})\leq D(X\parallel g_{X}),$
which implies the sharp lower bound
$h(Y)-h(X)\;\geq\;\frac{1}{2}\log\frac{\det(\Sigma_{X}+\sigma_{N}^{2}I)}{\det(\Sigma_{X})}.$ (9)
Equality holds if and only if $X$ is Gaussian.
Entropy–power inequality (EPI). Define the entropy power by
$\mathcal{E}(Z)=\frac{1}{2\pi e}\exp\!\Bigl{(}\frac{2}{n}h(Z)\Bigr{)}.$
Then the EPI states that for any independent $X$ and $N$,
$\mathcal{E}(X+N)\geq\mathcal{E}(X)+\mathcal{E}(N),$ (10)
which implies
$h(Y)-h(X)\;\geq\;\frac{n}{2}\log\!\Bigl{(}1+\frac{\mathcal{E}(N)}{\mathcal{E}(X)}\Bigr{)}>0.$ (11)
For $N\sim\mathcal{N}(0,\sigma_{N}^{2}I)$, we have $\mathcal{E}(N)=\sigma_{N}^{2}$. Since $D(Y\parallel g_{Y})\geq 0$, the decomposition Eq. (8) implies
$h(Y)-h(X)\;\leq\;\frac{1}{2}\log\frac{\det(\Sigma_{X}+\sigma_{N}^{2}I)}{\det(\Sigma_{X})}+D(X\parallel g_{X}).$ (12)
Equality holds if and only if $X$ is Gaussian.
Summary. Combining these facts, for arbitrary $X$ we obtain a sharp sandwich:
$\begin{split}&\frac{1}{2}\log\frac{\det(\Sigma_{X}+\sigma_{N}^{2}I)}{\det(\Sigma_{X})}\;\leq\;h(Y)-h(X)\\ &\leq\;\frac{1}{2}\log\frac{\det(\Sigma_{X}+\sigma_{N}^{2}I)}{\det(\Sigma_{X})}+D(X\parallel g_{X}).\end{split}$ (13)
The gap equals exactly the Gaussian value Eq. (6) if and only if $X$ is Gaussian. Building on this understanding, we can derive the following key insights and analyses:
@@ III-C Practical Estimation of Clean Image Entropy
Estimating the clean image entropy $H(I)$ in practice is non-trivial. We propose a deep neural network (DNN) to estimate $H(I)$ from the noisy observation $I_{n}$, an estimate of $H(I_{n})$, and (optionally) the noise variance $\sigma_{N}^{2}$. This leverages both the statistical properties of the noisy image and auxiliary entropy information.
Given a clean image $I\in\mathbb{R}^{H\times W\times C}$ and additive Gaussian noise $N\sim\mathcal{N}(0,\sigma_{N}^{2}I)$ with $I_{n}=I+N$, we first estimate:
@@
where $\hat{H}(I_{n})$ is obtained via mathematically lossless compression (MLLIC). Our DNN aims to approximate:
@@
with parameters $\phi$.
The network consists of two modules. First, a feature extractor $g_{\theta}(\cdot)$ with parameters $\theta$ processes $I_{n}$ through cascaded residual (ResBlocks) and downsampling residual blocks (Down-ResBlocks) to obtain:
@@
ResBlocks and Down-ResBlocks follow standard definitions with residual connections and optional downsampling to capture hierarchical features and noise characteristics.
The output feature $F(I_{n})$ is flattened into $F_{v}=\mathrm{vec}(F(I_{n}))$. We concatenate $F_{v}$, $\hat{H}(I_{n})$, and $\sigma_{N}$ into:
$Z=[F_{v};\hat{H}(I_{n});\sigma_{N}],$ (19)
which is passed to a regression module $r_{\psi}(\cdot)$ (an MLP with parameters $\psi$) to predict:
@@
We train the network by minimizing:
$\min_{\theta,\psi}\mathbb{E}[(\hat{H}(I)-H(I))^{2}],$ (21)
using datasets of $(I,I_{n},\sigma_{N})$ with known or high-fidelity $H(I)$.
In summary, this framework combines learned features from noisy inputs with entropy priors to effectively estimate clean image entropy under practical conditions.
@@ III-D Content-Adaptive Quantization and Probability Estimation
The proposed framework integrates entropy-guided adaptive quantization (see Figure 3) and probability estimation to optimize joint denoising and compression. By incorporating clean image entropy $\hat{H}(I)$, latent features $\mathbf{y}$, and hyperlatents $\mathbf{z}$, the framework achieves precise control over quantization and accurate probability modeling, effectively balancing rate-distortion trade-offs.
Figure 3: Illustration of the content-adaptive quantization process. Separate neural networks generate quantization step maps for the latent features $\mathbf{z}$ and hyperlatents $\mathbf{y}$, respectively.
@@
1. Quantization Map for Latent Features $\mathbf{y}$: The quantization step map $\mathbf{Q}_{\mathbf{y}}\in\mathbb{R}^{H\times W\times C}$ for the latent features $\mathbf{y}$ is computed as: $\mathbf{Q}_{\mathbf{y}}=f_{\mathbf{Q_{y}}}(\mathbf{y},\hat{\mathbf{z}},\hat{H}(I)),$ where $\mathbf{y}\in\mathbb{R}^{H\times W\times C}$ represents the latent features, $\hat{\mathbf{z}}$ denotes the decoded hyperlatents, and $\hat{H}(I)$ is the estimated clean image entropy. The function $f_{\mathbf{Qy}}(\cdot)$ is a neural network parameterized to produce $\mathbf{Q}_{\mathbf{y}}$ by analyzing the spatial, channel-wise, and entropy-based characteristics of $\mathbf{y}$.
The quantization process for $\mathbf{y}$ is defined as: $\mathbf{y}_{q}=\mathrm{round}\left(\mathbf{y}/\mathbf{Q}_{\mathbf{y}}\right),$ where division and rounding are performed element-wise. Regions with higher complexity or structural importance are assigned finer quantization steps, ensuring accurate reconstruction, while smoother regions are quantized more coarsely to save bits.
2. Quantization Map for Hyperlatents $\mathbf{z}$: The quantization step map $\mathbf{Q}_{\mathbf{z}}\in\mathbb{R}^{H^{\prime}\times W^{\prime}\times C^{\prime}}$ for the hyperlatents $\mathbf{z}$ is derived as: $\mathbf{Q}_{\mathbf{z}}=f_{\mathbf{Q_{z}}}(\mathbf{z},\hat{H}(I)),$ where $\mathbf{z}\in\mathbb{R}^{H^{\prime}\times W^{\prime}\times C^{\prime}}$ represents the hyperlatents, and $f_{\mathbf{Qz}}(\cdot)$ is a neural network conditioned on $\mathbf{z}$ and $\hat{H}(I)$. This design ensures that the quantization map adapts to the global statistical properties of $\mathbf{z}$ and the entropy of the clean image.
The quantization process for $\mathbf{z}$ follows: $\mathbf{z}_{q}=\mathrm{round}\left(\mathbf{z}/\mathbf{Q}_{\mathbf{z}}\right).$ This adaptive quantization efficiently encodes the hyperlatents while minimizing distortion.